# Calibrating LLM Confidence by Probing Perturbed Representation Stability

### v1

@@ 6 Discussion
The CCPS Framework Provides an Efficient and Scalable Approach to Confidence Estimation.  CCPS is designed to be lightweight. Once features are extracted, the confidence estimation model itself consists of relatively small MLPs or CNNs (as detailed in Appendix E), making its training and inference efficient. This contrasts with methods like CT which, despite using LoRA, require fine-tuning larger portions of the base LLM and can be resource-intensive (e.g., CT reportedly takes $\sim$4 GPU days on an NVIDIA V100). Furthermore, CCPS avoids some scalability concerns present in other methods. For example, LitCab’s projection layer size (hidden_dim $\times$ vocabulary_size) can become very large for LLMs with extensive vocabularies, and its reliance on multiple negative samples per question for its contrastive learning imposes specific data curation requirements. CCPS, on the other hand, uses more compact projection networks and only requires labels of correctness for the LLM’s generated answers.
Probing Internal Representational Stability Forms the Core of CCPS’s Mechanism.  The methodological foundation of CCPS lies in quantifying internal consistency. While prior work has highlighted output consistency (e.g., through multiple sampling of responses) as a strong indicator of LLM reliability Zhou et al. (2025), such approaches can be computationally expensive. CCPS internalizes this concept by perturbing hidden state representations. The premise is that if an LLM is truly confident, its internal decision-making process for a token should be stable against relevant adversarial nudges. Our results suggest that features derived from this internal stability analysis serve as effective proxies for confidence.
@@
Perturbation-Derived Features Offer Key Insights into LLM Confidence Signals.  The SHAP value analyses (Appendix I.5) provide insights into which features derived from our perturbation process are most influential. Consistently across different LLMs and datasets, the original entropy of the LLM’s output distribution for a token emerges as an important feature. As expected, higher original entropy typically shows a negative correlation with the prediction of correctness (meaning higher entropy contributes to predicting the answer as incorrect), signifying that greater initial uncertainty in the LLM’s choice is indicative of a potentially incorrect answer. More revealingly, many of the top-ranking features are those derived from the perturbed states. For instance, the margin between the logits of the top-ranked and second-ranked tokens after perturbation often shows a positive correlation with correctness; a larger margin, even under adversarial stress, indicates a more decisive and less ambiguous output from the LLM, which CCPS learns as a sign of confidence. Similarly, a higher epsilon-to-flip value, indicating that a larger perturbation magnitude is needed to make the LLM change its predicted token, consistently contributes positively to the confidence score. These findings affirm that the dynamic response to perturbation, not just the initial state, provides critical signals for confidence estimation.
@@ Acknowledgments
This work was supported by the Henry Ford Health System and JP Morgan Chase Faculty Research Award. The authors are solely responsible for the contents of the paper and the opinions expressed in this publication do not reflect those of the funding agencies. The authors also acknowledge the use of Large Language Models to assist in polishing the language and grammar of this manuscript.
@@ Appendix B Feature Set Description
This appendix details the features extracted for analyzing the language model’s token-level generative behavior. These features, defined in Table 2, are derived from the model’s internal states and its responses to systematic perturbations. The perturbations are applied to a token’s hidden state by incrementally moving it along the normalized Jacobian vector of the actual observed token’s log-probability, across a predefined number of steps and range of magnitudes.
@@ B.3 Perturbed State Features
[…]
@@
[…]
@@
original_entropy Entropy of the original predictive distribution: $-\sum_{i}P_{\text{original}}(i)\log P_{\text{original}}(i)$.
@@
pei_value_token Perturbation-Effect Integration (PEI): total normalized drop in log-prob of the actual token over all perturbation steps.
@@
[…]
@@ C.1 Training and Validation Datasets
For training and validating our confidence estimation models, we utilize the CT-CHOICE and CT-OE datasets, designed for multiple-choice and open-ended question-answering formats, respectively. These datasets, generated following the methodology of Kapoor et al. (2024b), aggregate a diverse collection of commonly used public question-answering datasets. Instances from these datasets were formatted to ensure a maximum input sequence length of 1,600 tokens during our training process. The underlying datasets include:
@@ C.3 Response Generation and Labeling
For all datasets described above, responses from the base LLMs were first generated to create the instances for our confidence estimation task. The user prompt, which includes the question and any contextual information (such as few-shot exemplars), was constructed following the methodology of Kapoor et al. (2024b), to which we refer the reader for further details. We employed specific system prompts for guiding the base LLMs during response generation, as detailed in Table 5. These prompts are similar to those used by Kapoor et al. (2024b) but were slightly refined for improved clarity to the LLMs. In line with their approach, for multiple-choice question-answering datasets (CT-CHOICE, MMLU-CHOICE, MMLU-PRO-CHOICE), answers were generated with a maximum token limit of 1, corresponding to the chosen option letter. For open-ended datasets (CT-OE, MMLU-OE), responses were generated using greedy decoding with a maximum length of 30 tokens.
@@
Each generated response was subsequently labeled as correct or incorrect. For multiple-choice questions, correctness was determined by a straightforward string match between the LLM’s generated option letter and the ground truth option. For open-ended responses, assessing semantic equivalence requires a more nuanced approach. Kapoor et al. (2024b) conducted a comparative analysis of different grading techniques against human evaluations. Their study found that GPT-4 assessments exhibited a low average absolute difference of 4.5% in accuracy estimation compared to human annotators, while GPT-3.5 Turbo also demonstrated strong agreement, with an average difference of 8.7%. Despite the superior performance of GPT-4, they ultimately employed GPT-3.5 Turbo for their labeling due to expediency. Leveraging their validation of advanced LLMs for this task, and given the availability of even more capable models since their study, we utilized the more recent GPT-4o-mini model via its API for assessing the correctness of open-ended LLM responses. This choice was made to ensure the highest quality semantic equivalence judgments. The prompts used for this grading task are detailed in Table 6.
@@ G.4 Discussion of Parameter Efficiency
Meta-Llama-3.1-8B-Instruct 55,069×55,069$\times$55 , 069 × 440×440$\times$440 ×
Qwen2.5-14B-Instruct 81,610×81,610$\times$81 , 610 × 824×824$\times$824 ×
Mistral-Small-24B-Instruct-2501 70,344×70,344$\times$70 , 344 × 687×687$\times$687 ×
Qwen2.5-32B-Instruct 81,610×81,610$\times$81 , 610 × 1,099×1,099$\times$1 , 099 ×
@@ H.2 Brier Score
Brier $\text{Brier Score}=\frac{1}{N}\sum_{k=1}^{N}(p_{k}-o_{k})^{2}$
@@ H.5 Area Under the Receiver Operating Characteristic Curve (AUROC)
Appendix I Extended Results and Analyses
@@
I.1 Per-Dataset Aggregate Performance Tables
To offer a comprehensive comparison of all evaluated methods, including all baselines, Tables 12, 13, and 14 present aggregate performance metrics for the MMLU-CHOICE, MMLU-PRO-CHOICE, and MMLU-OE datasets, respectively. Unlike the main paper’s Table 1 which shows mean scores across tasks for selected methods, these tables detail the mean $\pm$ standard deviation for all methods across all evaluated LLMs for each metric, providing insight into the consistency of performance.
@@ I.1 Per-Dataset Aggregate Performance Tables
Meta-Llama-3.1-8B-Instruct P(True) 35.9$\pm$5.7 39.4$\pm$4.2 45.4$\pm$5.9 66.0$\pm$14.6 49.2$\pm$5.7
P(IK) 18.9$\pm$9.6 25.4$\pm$2.4 63.9$\pm$14.8 65.3$\pm$14.9 49.8$\pm$1.8
LTS 28.9$\pm$6.6 34.5$\pm$3.7 44.6$\pm$6.9 66.6$\pm$14.1 50.1$\pm$4.3
IT 33.4$\pm$5.3 37.5$\pm$3.7 47.2$\pm$4.9 66.5$\pm$14.7 49.8$\pm$5.0
SAPLMA-M 17.9$\pm$9.7 24.8$\pm$2.7 64.9$\pm$15.1 64.9$\pm$15.3 49.5$\pm$3.1
SAPLMA-UM 18.1$\pm$9.8 24.9$\pm$2.7 64.9$\pm$15.1 64.6$\pm$15.4 49.3$\pm$3.3
SAPLMA-F 18.2$\pm$9.8 24.9$\pm$2.6 64.9$\pm$15.0 65.0$\pm$15.0 49.6$\pm$2.3
CT 10.7$\pm$6.7 21.1$\pm$5.7 67.8$\pm$12.2 74.2$\pm$15.5 62.8$\pm$8.0
LitCab 10.9$\pm$4.8 18.1$\pm$5.5 73.2$\pm$8.7 84.0$\pm$13.5 77.1$\pm$8.2
CCPS 6.5$\pm$3.9 17.1$\pm$4.7 73.4$\pm$8.5 84.1$\pm$13.5 77.1$\pm$8.5
Qwen2.5-14B-Instruct P(True) 47.0$\pm$6.2 47.0$\pm$4.8 41.3$\pm$6.4 79.2$\pm$12.5 51.2$\pm$5.8
P(IK) 25.1$\pm$13.0 24.1$\pm$3.1 76.8$\pm$12.2 78.3$\pm$12.1 49.9$\pm$2.4
LTS 41.5$\pm$6.5 43.0$\pm$4.3 38.6$\pm$6.2 78.9$\pm$12.6 49.7$\pm$5.7
IT 44.7$\pm$5.9 44.0$\pm$5.2 45.7$\pm$7.1 79.4$\pm$12.4 50.4$\pm$6.7
SAPLMA-M 23.8$\pm$13.2 23.0$\pm$4.0 78.1$\pm$12.1 78.4$\pm$12.3 50.5$\pm$3.0
SAPLMA-UM 23.7$\pm$13.2 23.0$\pm$4.0 78.2$\pm$12.1 78.4$\pm$12.1 50.3$\pm$2.4
SAPLMA-F 24.0$\pm$12.9 23.0$\pm$3.7 78.1$\pm$12.1 78.5$\pm$12.2 50.3$\pm$3.0
CT 12.1$\pm$8.1 17.0$\pm$8.1 78.6$\pm$11.5 84.7$\pm$10.9 64.8$\pm$9.1
LitCab 45.6$\pm$11.3 20.0$\pm$10.8 78.3$\pm$12.0 83.7$\pm$10.2 65.3$\pm$5.4
CCPS 6.3$\pm$3.7 13.1$\pm$5.8 80.2$\pm$9.5 92.1$\pm$8.1 81.6$\pm$7.0
Mistral-Small-24B-Instruct-2501 P(True) 42.1$\pm$8.5 43.3$\pm$5.7 38.1$\pm$7.9 80.5$\pm$12.1 49.3$\pm$8.0
P(IK) 12.4$\pm$9.3 17.8$\pm$8.7 73.9$\pm$18.8 82.6$\pm$12.2 56.3$\pm$8.9
LTS 36.2$\pm$9.0 38.3$\pm$4.7 36.1$\pm$8.4 80.2$\pm$12.6 49.2$\pm$6.2
IT 37.3$\pm$7.3 39.4$\pm$5.0 42.9$\pm$7.7 81.3$\pm$12.0 49.8$\pm$7.9
SAPLMA-M 24.5$\pm$14.0 22.5$\pm$4.0 79.8$\pm$12.9 79.9$\pm$12.8 49.8$\pm$2.0
SAPLMA-UM 24.6$\pm$14.1 22.5$\pm$4.1 79.8$\pm$12.9 80.1$\pm$12.9 50.6$\pm$2.9
SAPLMA-F 25.2$\pm$14.3 22.9$\pm$4.1 79.8$\pm$12.9 79.8$\pm$12.9 49.8$\pm$2.3
CT 8.2$\pm$7.4 15.5$\pm$7.8 79.6$\pm$13.1 83.3$\pm$11.5 56.5$\pm$7.6
LitCab 13.5$\pm$6.7 15.1$\pm$7.4 79.5$\pm$9.8 91.5$\pm$8.4 78.2$\pm$8.0
CCPS 5.8$\pm$3.2 11.5$\pm$6.0 83.0$\pm$10.3 93.1$\pm$7.8 83.3$\pm$7.6
Qwen2.5-32B-Instruct P(True) 44.0$\pm$7.0 45.7$\pm$5.5 41.9$\pm$7.4 84.0$\pm$10.3 52.1$\pm$7.3
P(IK) 28.6$\pm$12.7 23.5$\pm$4.2 81.7$\pm$10.7 82.6$\pm$10.4 49.9$\pm$2.9
LTS 37.1$\pm$6.7 40.2$\pm$4.4 41.9$\pm$7.4 84.1$\pm$10.3 52.2$\pm$7.3
IT 41.9$\pm$7.6 44.0$\pm$6.2 44.6$\pm$8.3 84.6$\pm$10.5 54.8$\pm$7.6
SAPLMA-M 27.3$\pm$13.2 22.7$\pm$4.7 82.3$\pm$10.6 82.4$\pm$10.6 49.7$\pm$4.2
SAPLMA-UM 27.7$\pm$12.8 22.8$\pm$4.7 82.3$\pm$10.6 82.3$\pm$10.7 49.4$\pm$3.6
SAPLMA-F 27.2$\pm$12.9 22.5$\pm$4.5 82.3$\pm$10.6 82.4$\pm$10.7 49.9$\pm$2.8
CT 45.2$\pm$7.0 46.9$\pm$5.1 37.2$\pm$6.1 84.3$\pm$10.1 51.6$\pm$8.0
LitCab 43.2$\pm$11.0 15.9$\pm$9.3 82.6$\pm$10.4 87.9$\pm$7.9 67.2$\pm$6.5
CCPS 6.3$\pm$3.1 10.8$\pm$5.2 84.1$\pm$8.9 94.1$\pm$5.9 82.8$\pm$6.9
@@
Meta-Llama-3.1-8B-Instruct P(True) 25.3$\pm$6.7 33.1$\pm$4.5 54.8$\pm$6.8 37.1$\pm$11.7 49.8$\pm$2.1
P(IK) 41.7$\pm$15.6 44.1$\pm$11.5 38.2$\pm$11.7 37.3$\pm$13.7 49.9$\pm$3.2
LTS 17.0$\pm$6.7 29.1$\pm$3.5 55.4$\pm$7.0 36.9$\pm$12.3 49.9$\pm$2.2
IT 26.8$\pm$4.7 33.8$\pm$2.9 52.8$\pm$4.8 37.7$\pm$12.2 50.0$\pm$2.7
SAPLMA-M 40.4$\pm$14.0 40.3$\pm$8.6 36.7$\pm$12.7 37.3$\pm$13.0 50.1$\pm$1.8
SAPLMA-UM 41.0$\pm$14.3 41.0$\pm$9.1 36.7$\pm$12.7 37.5$\pm$13.3 50.3$\pm$3.0
SAPLMA-F 40.2$\pm$14.9 40.7$\pm$10.0 36.8$\pm$12.7 37.2$\pm$12.9 50.3$\pm$1.8
CT 21.5$\pm$11.5 29.8$\pm$5.9 50.4$\pm$11.7 43.7$\pm$14.4 57.3$\pm$4.4
LitCab 16.6$\pm$2.9 24.7$\pm$2.6 66.1$\pm$4.2 51.7$\pm$18.4 63.6$\pm$9.0
CCPS 4.5$\pm$2.1 20.0$\pm$2.2 70.4$\pm$4.0 55.2$\pm$19.4 67.9$\pm$8.1
Qwen2.5-14B-Instruct P(True) 33.7$\pm$7.1 38.6$\pm$4.8 49.9$\pm$6.0 55.4$\pm$12.6 51.4$\pm$1.3
P(IK) 27.3$\pm$11.4 33.9$\pm$8.2 53.6$\pm$11.5 53.5$\pm$13.4 49.1$\pm$2.3
LTS 26.7$\pm$7.4 34.5$\pm$4.5 49.3$\pm$6.7 54.6$\pm$11.8 50.7$\pm$2.2
IT 33.5$\pm$6.0 38.3$\pm$3.6 50.5$\pm$4.2 55.6$\pm$12.1 51.1$\pm$2.4
SAPLMA-M 28.1$\pm$13.2 33.4$\pm$8.6 53.4$\pm$12.5 54.1$\pm$13.3 50.1$\pm$3.0
SAPLMA-UM 27.4$\pm$13.5 33.0$\pm$8.6 53.5$\pm$12.5 53.8$\pm$13.0 49.9$\pm$3.1
SAPLMA-F 25.7$\pm$12.3 32.1$\pm$7.7 53.4$\pm$12.5 53.8$\pm$12.5 49.3$\pm$2.8
CT 20.4$\pm$10.3 28.7$\pm$6.3 55.6$\pm$11.4 59.4$\pm$12.9 56.6$\pm$3.5
LitCab 49.7$\pm$4.2 38.3$\pm$8.8 55.3$\pm$11.6 66.2$\pm$10.1 68.0$\pm$3.7
CCPS 4.2$\pm$1.8 20.1$\pm$2.9 69.2$\pm$5.4 75.8$\pm$10.5 74.0$\pm$4.8
Mistral-Small-24B-Instruct-2501 P(True) 32.0$\pm$8.1 37.2$\pm$5.0 46.9$\pm$7.3 57.5$\pm$12.3 50.2$\pm$2.2
P(IK) 32.3$\pm$11.4 36.3$\pm$9.7 56.1$\pm$10.9 57.4$\pm$13.4 50.6$\pm$2.1
LTS 24.7$\pm$7.6 32.7$\pm$3.7 46.2$\pm$7.2 56.6$\pm$12.3 49.2$\pm$1.7
IT 31.2$\pm$6.7 36.2$\pm$3.9 47.0$\pm$6.1 58.4$\pm$12.0 50.3$\pm$2.8
SAPLMA-M 24.5$\pm$12.1 30.7$\pm$8.0 56.7$\pm$12.4 57.0$\pm$13.3 49.9$\pm$2.8
SAPLMA-UM 24.5$\pm$11.7 30.7$\pm$8.0 56.7$\pm$12.4 57.6$\pm$13.4 50.6$\pm$3.1
SAPLMA-F 25.1$\pm$12.8 31.4$\pm$8.5 56.7$\pm$12.4 56.8$\pm$12.2 49.8$\pm$2.2
CT 17.8$\pm$9.7 27.4$\pm$5.9 58.2$\pm$11.6 60.1$\pm$13.1 54.3$\pm$3.1
LitCab 32.2$\pm$3.1 34.6$\pm$3.2 57.0$\pm$3.7 66.2$\pm$12.8 60.1$\pm$5.0
CCPS 4.5$\pm$1.9 18.6$\pm$3.3 71.3$\pm$6.4 79.5$\pm$9.4 77.2$\pm$5.2
Qwen2.5-32B-Instruct P(True) 34.6$\pm$6.8 39.5$\pm$4.9 46.1$\pm$5.9 60.1$\pm$12.2 50.3$\pm$2.7
P(IK) 23.6$\pm$9.9 30.8$\pm$7.7 58.0$\pm$10.8 59.5$\pm$11.9 50.2$\pm$2.5
LTS 27.5$\pm$6.7 34.8$\pm$3.9 46.1$\pm$5.9 60.1$\pm$12.2 50.3$\pm$2.7
IT 36.6$\pm$6.7 40.9$\pm$5.3 45.9$\pm$5.9 60.1$\pm$12.1 51.0$\pm$2.7
SAPLMA-M 24.8$\pm$12.0 30.5$\pm$8.3 59.3$\pm$11.8 59.9$\pm$12.1 49.9$\pm$2.8
SAPLMA-UM 26.9$\pm$12.1 31.8$\pm$8.9 59.3$\pm$11.8 60.2$\pm$12.0 49.8$\pm$3.3
SAPLMA-F 23.7$\pm$11.2 30.0$\pm$7.9 59.3$\pm$11.8 59.4$\pm$11.6 49.5$\pm$2.7
CT 38.0$\pm$8.5 41.6$\pm$6.4 44.8$\pm$7.2 60.5$\pm$11.3 49.9$\pm$2.7
LitCab 48.4$\pm$3.5 33.7$\pm$8.7 60.8$\pm$11.0 72.7$\pm$8.8 70.3$\pm$4.7
CCPS 4.6$\pm$2.1 18.5$\pm$3.4 71.8$\pm$6.1 82.4$\pm$7.7 77.8$\pm$4.7

### v2

@@ 6 Discussion
The CCPS Framework Provides an Efficient and Scalable Approach to Confidence Estimation.  CCPS is designed to be lightweight. Once features are extracted, the confidence estimation model itself consists of relatively small MLPs or CNNs (as detailed in Appendix E), making its training and inference efficient. Specifically for our OE models, the convolutional architecture proves to be an effective design choice; a Token Masking Impact Analysis (detailed in Appendix K.1) demonstrates that the model successfully learns to prioritize semantically meaningful tokens by being significantly more sensitive to content words than to grammatical filler words, all within a compact architecture. This efficiency contrasts sharply with methods like CT, which, despite using LoRA, require fine-tuning larger portions of the base LLM and can be resource-intensive (e.g., CT reportedly takes $\sim$4 GPU days on an NVIDIA V100). Furthermore, CCPS avoids some scalability concerns present in other methods. For example, LitCab’s projection layer size (hidden_dim $\times$ vocabulary_size) can become very large for LLMs with extensive vocabularies, and its reliance on multiple negative samples per question for its contrastive learning imposes specific data curation requirements. CCPS, on the other hand, uses more compact projection networks and only requires labels of correctness for the LLM’s generated answers.
Probing Internal Representational Stability Forms the Core of CCPS’s Mechanism. The methodological foundation of CCPS lies in quantifying internal consistency. Prior work has shown that external output consistency is a useful reliability signal—for example, generating multiple responses and measuring consensus, as in the Self-Consistency (SC) method Xiong et al. (2024). Such approaches, however, are computationally costly because they require repeated full-generation passes Zhou et al. (2025). CCPS internalizes this idea by perturbing hidden state representations instead. The premise is that if an LLM is truly confident, its internal decision-making process for a token should remain stable under small, targeted perturbations. Our results suggest that features derived from this stability serve as effective proxies for confidence. Furthermore, our direct empirical comparison in Appendix K.2 shows that CCPS consistently outperforms SC, particularly in calibration, validating our approach as a more efficient and effective way to measure consistency. To further isolate the source of these gains, Appendix K.3 presents a detailed ablation study contrasting feature sets from unperturbed states, perturbation-derived features, and their combination. The results confirm that perturbation features are the dominant driver of CCPS’s strong performance, especially on more difficult tasks.
@@
Perturbation-Derived Features Offer Key Insights into LLM Confidence Signals.  The SHAP value analyses (Appendix K.4 provide insights into which features derived from our perturbation process are most influential. Consistently across different LLMs and datasets, the original entropy of the LLM’s output distribution for a token emerges as an important feature. As expected, higher original entropy typically shows a negative correlation with the prediction of correctness (meaning higher entropy contributes to predicting the answer as incorrect), signifying that greater initial uncertainty in the LLM’s choice is indicative of a potentially incorrect answer. More revealingly, many of the top-ranking features are those derived from the perturbed states. For instance, the margin between the logits of the top-ranked and second-ranked tokens after perturbation often shows a positive correlation with correctness; a larger margin, even under adversarial stress, indicates a more decisive and less ambiguous output from the LLM, which CCPS learns as a sign of confidence. Similarly, a higher epsilon-to-flip value, indicating that a larger perturbation magnitude is needed to make the LLM change its predicted token, consistently contributes positively to the confidence score. These findings affirm that the dynamic response to perturbation, not just the initial state, provides critical signals for confidence estimation. To further validate these findings and rigorously quantify the importance of each feature, we conducted a comprehensive leave-one-out ablation study, detailed in Appendix K.5.
@@ Acknowledgments
This work was supported in part by the Henry Ford Health + Michigan State University Health Sciences Cancer Seed Funding Program and by the JPMorgan Chase AI Research Faculty Research Award. The authors are solely responsible for the contents of this paper; the opinions expressed do not necessarily reflect those of the funding organizations. The authors also acknowledge the use of Large Language Models to assist in polishing the language and grammar of this manuscript.
@@ Appendix B Feature Set Description
This appendix details the features extracted for analyzing the language model’s token-level generative behavior. Our final model uses a 75-dimensional feature vector ($D_{f}=75$) per token, which is derived from 30 base metric definitions. These features are calculated from the model’s internal states and its responses to systematic perturbations, where a token’s hidden state is incrementally moved along an adversarial direction. The 75 features are composed of 15 single scalar values (12 Original State features and 3 Overall Perturbation features), and 60 features derived from statistical summaries (mean, standard deviation, minimum, and maximum) of the remaining 15 base metrics that are computed across the perturbation trajectory. This process yields a total of $15+(15\times 4)=75$ features per token. The definitions for the base metrics are provided in Table 2.
@@ B.3 Perturbed State Features
original_log_prob_actual Log-probability of the actual token based on the model’s original (unperturbed) output distribution, i.e. $\log P_{\text{original}}(\text{actual\_token})$.
original_prob_actual Probability of the actual token based on the model’s original output distribution, i.e. $P_{\text{original}}(\text{actual\_token})$.
@@
original_prob_argmax Highest probability assigned to any token by the original model, i.e. $P_{\text{original}}(\text{argmax\_token})$.
@@
original_entropy Entropy of the original predictive distribution: $\sum_{i}P_{\text{original}}(i)\log P_{\text{original}}(i)$.
@@
pei_value_token Perturbation Energy Integral (PEI): total normalized drop in log-prob of the actual token over all perturbation steps.
@@
perturbed_log_prob_actual Log-prob of the actual token after hidden-state perturbation, $\log P_{\text{perturbed}}(\text{actual\_token})$.
perturbed_prob_actual Probability of the actual token after perturbation, $P_{\text{perturbed}}(\text{actual\_token})$.
@@ C.1 Training and Validation Datasets
For training and validating our confidence estimation models, we utilize the CT-CHOICE and CT-OE datasets, designed for multiple-choice and open-ended QA formats, respectively. These datasets, generated following the methodology of Kapoor et al. (2024b), aggregate a diverse collection of commonly used public QA datasets. Instances from these datasets were formatted to ensure a maximum input sequence length of 1,600 tokens during our training process. The underlying datasets include:
@@ C.3 Response Generation and Labeling
For all datasets described above, responses from the base LLMs were first generated to create the instances for our confidence estimation task. The user prompt, which includes the question and any contextual information (such as few-shot exemplars), was constructed following the methodology of Kapoor et al. (2024b), to which we refer the reader for further details. We employed specific system prompts for guiding the base LLMs during response generation, as detailed in Table 5. These prompts are similar to those used by Kapoor et al. (2024b) but were slightly refined for improved clarity to the LLMs. In line with their approach, for multiple-choice QA datasets (CT-CHOICE, MMLU-CHOICE, MMLU-PRO-CHOICE), answers were generated with a maximum token limit of 1, corresponding to the chosen option letter. For open-ended datasets (CT-OE, MMLU-OE), responses were generated using greedy decoding with a maximum length of 30 tokens.
@@
Each generated response was subsequently labeled as correct or incorrect. For multiple-choice questions, correctness was determined by a straightforward string match between the LLM’s generated option letter and the ground truth option. For open-ended responses, assessing semantic equivalence requires a more nuanced approach. To this end, and consistent with recent literature, we employed a powerful auxiliary LLM not as a knowledge oracle, but as a semantic equivalence assessor. The grader model’s task was constrained: for each question, it was provided with the ground-truth answer available in the dataset and the LLM’s generated response, and was prompted to determine if the two answers were semantically equivalent. The reliability of using a powerful LLM for this specific equivalence task has been validated by Kapoor et al. (2024b), who conducted a comparative analysis against human evaluations. Their study found that GPT-4 assessments exhibited a low average absolute difference of 4.5% in accuracy estimation compared to human annotators. Building upon their findings, and given the availability of even more capable models since their study, we utilized the more recent GPT-4o-mini model to ensure the highest quality semantic equivalence judgments. The prompts used for this grading task are detailed in Table 6.
@@ G.4 Discussion of Parameter Efficiency
Meta-Llama-3.1-8B-Instruct $55,069\times$ $440\times$
Qwen2.5-14B-Instruct $81,610\times$ $824\times$
Mistral-Small-24B-Instruct-2501 $70,344\times$ $687\times$
Qwen2.5-32B-Instruct $81,610\times$ $1,099\times$
@@ H.2 Brier Score
$\text{Brier Score}=\frac{1}{N}\sum_{k=1}^{N}(p_{k}-o_{k})^{2}$
@@ H.5 Area Under the Receiver Operating Characteristic Curve (AUROC)
Appendix I Evaluation on a High-Stakes Domain: MedMCQA
To further assess the cross-domain robustness of CCPS and validate its performance in a critical, high-stakes setting as motivated in our introduction, we conducted an additional set of experiments on the MedMCQA benchmark Pal et al. (2022). This medical QA dataset provides a specialized domain to test the generalizability of our method. We compared CCPS against the strong Calibration Tuning (CT) baseline on all four base LLMs, using the identical experimental settings and fixed hyperparameters from our main evaluations to ensure a fair comparison. The performance on both multiple-choice and open-ended formats is presented in Table 12.
Dataset Model Method ECE $\downarrow$ Brier $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
MedMCQA (Multiple-Choice) Meta-Llama-3.1-8B CT 26.8 31.7 51.1 56.4 54.7
CCPS 11.0 20.9 67.7 74.0 73.9
Qwen2.5-14B CT 28.0 30.1 62.2 68.2 55.3
CCPS 13.8 20.5 66.1 80.0 73.9
Mistral-Small-24B CT 23.4 27.9 64.3 69.4 53.7
CCPS 21.2 25.4 65.2 70.8 54.5
Qwen2.5-32B CT 41.1 41.8 44.0 72.4 52.4
CCPS 12.7 19.0 70.2 87.2 76.8
MedMCQA (Open-Ended) Meta-Llama-3.1-8B CT 12.1 18.7 71.8 38.7 66.4
CCPS 9.1 18.9 74.3 45.0 74.7
Qwen2.5-14B CT 14.0 22.2 59.3 41.2 65.5
CCPS 11.5 21.2 68.5 44.5 68.6
Mistral-Small-24B CT 14.6 29.7 63.4 36.6 53.9
CCPS 12.4 22.3 69.3 45.0 62.4
Qwen2.5-32B CT 18.5 24.6 67.0 31.8 56.2
CCPS 18.3 21.9 72.4 43.5 70.1
Table 12: Performance on MedMCQA across two setups: Multiple-Choice and Open-Ended. Arrows indicate whether lower ($\downarrow$) or higher ($\uparrow$) values are better. Best results per model are bolded.
The results on this specialized medical dataset show that CCPS’s performance advantages are not only maintained but often amplified. In the multiple-choice setting, CCPS substantially improves upon CT across all metrics, reducing ECE by over 59% for Llama-8B and nearly 69% for Qwen-32B. In the open-ended setting, CCPS again demonstrates superior calibration and discrimination across nearly all models and metrics. These new results provide strong evidence that the advantages of CCPS generalize robustly beyond standard knowledge benchmarks to this critical, high-stakes domain.
Appendix J Extended Results and Analyses
@@ Appendix I Evaluation on a High-Stakes Domain: MedMCQA
J.1 Per-Dataset Aggregate Performance Tables
To offer a comprehensive comparison of all evaluated methods, including all baselines, Tables 13, 14, and 15 present aggregate performance metrics for the MMLU-CHOICE, MMLU-PRO-CHOICE, and MMLU-OE datasets, respectively. Unlike the main paper’s Table 1 which shows mean scores across tasks for selected methods, these tables detail the mean $\pm$ standard deviation for all methods across all evaluated LLMs for each metric, providing insight into the consistency of performance.
@@
[…]
@@
[…]